# biblio.bib Resolution Limiting Factors in Low-Energy Cascade Zenith Reconstruction with the IceCube Upgrade

### v1

@@ 4.1 Reconstruction techniques in IceCube
A point-like cascade event is characterized by 6 parameters [Abbasi:2022]: the incoming zenith and azimuth of the shower axis, interaction vertex coordinates = {vx,vy,vz} and the vertex time tv. The average number of emitted photons scales linearly with the summed track length of all charged particles, which increases linearly with the energy of the particle [RADEL2013102] [Wiebusch:1995bw]. This results in a sparse pulse distribution, with very few detected photons per event at energies of a few GeV [Abbasi:2022]. Noise cleaning further reduces the physics hits due to misclassification posing challenges to the reconstruction process and necessitating an accurate algorithm.
An IceCube reconstruction algorithm employed in the low-energy regime is the Single-string ANTares-inspired Analysis (SANTA) [Abbasi:2022] which utilizes information from minimally scattered photons. This is done by a rigorous photon hit selection based on the delay timing of photons followed by an event selection based on the surviving photon hit counts. A chi-squared fit on the measured arrival photon hit times for an event hypothesis is finally performed to obtain the optimized parameters. Another standard reconstruction, which additionally exploits the in-ice scattering information, is a maximum likelihood estimation [Stigler:2007] that fits a point-like cascade model to the measured light yield in the detector. The cascade template outlines the expected light distribution in a module for the given cascade hypothesis parameters. A binned Poisson likelihood finally identifies the hypothesis that best describes the observed light pattern.
The IceCube Upgrade will feature new detector strings spaced irregularly and novel optical modules with different PMTs per module. Adapting the likelihood methods to accommodate these upgrades is challenging because incorporating them breaks the symmetries in the current detector such as its regular geometry and uniform module type, which makes it hard to keep memory usage and computing time manageable. More recent algorithms use machine learning models such as Convolutional Neural Networks (CNNs) [Jessie_Micalf] and Graph Neural Networks (GNNs) [GNN] trained on high-statistics, low-energy data for atmospheric neutrino oscillation studies [Jessie_Micalf]. The CNNs transform the hit module map into uniformly sized images that assume a regular detector geometry and significantly reduce the event information [GNN], whereas, the GNNs incorporate the irregular geometry naturally through a graph representation of the data. The GNN architecture also integrates the PMT pointing directions and different module types which increases the parameter space but remains manageable due to the fast reconstruction speeds. Overall, they have demonstrated a 15-20% improvement [GNN] in directional resolution compared to the traditional maximum likelihood methods used in IceCube [Abbasi:2022].
@@ 4.2 Reconstruction framework for this study
The longitudinal length of hadronic showers in water and ice increases logarithmically with the shower energy and measures 2-4 meters at energies of O(10 GeV) [Steuer_theses] [R_del:2013] [shower_KM3NeT]. Since the sparse instrumentation of neutrino telescopes does not allow the resolution of shower details at such low energies, we model it as a point-like source with an anisotropic photon emission. The parameter of interest for this study is the zenith angle , derived from reconstructing the direction unit vectors = {, , } of the shower axis in this study, is a relevant parameter for oscillation analyses. Combined with the parameters listed in Section 4.1, this study aims to reconstruct 7 parameters - - sufficient to describe this topology.
The reconstruction uses an extended unbinned likelihood approach for its statistical robustness, with the assumption that it achieves minimum variance bounds since the assumed PDFs accurately represent the true distributions. Our method uses the direction, residual timing, and charge information of the detected photons, followed by a combined likelihood that exploits the correlation between these observables which is finally used for reconstructions in Section 5. For an anisotropic point source, the angular distribution of the incident photons relative to the direction of the shower axis encloses directional information about the shower axis because more photons are detected in the modules in front of the shower. For the same reason, the number of photons registered by sensors is direction-dependent because sensors in front of a shower receive more photons than those behind it. The residual timings, defined as the difference between the observed and geometric timestamps, are sensitive to shower direction since sensors behind the shower receive more backscattered and delayed photons.
@@
where is the 7-dimensional parameter set of the event, is the module index, is the photon index for a given module index , and N denotes the total number of modules. The opening angle enclosed between the shower axis and a detected photon is represented by , and represents the angle between the shower axis and a vector connecting the vertex and a module. The residual timing of a detected photon, defined as the difference between observed and geometric timestamps, is denoted by . The photon direction, timing, and correlation PDFs, , , and are the expected averaged distributions over all hit sensors and are not individually defined for each sensor. The photon count per module, or charge, follows a Poissonian distribution given by
@@
Equation 4.1 calculates the likelihood of observing a given angular distribution of detected photons and hit modules relative to the shower axis for different parameter hypotheses (see left panel of Fig. 3). This formulation exploits the radial symmetry of the photon emission relative to the shower axis to infer its direction. When module resolutions are considered, the photon direction is replaced by the axis direction of the hit PMT (see Fig. 3). Equation 4.2 evaluates the likelihood based on the observed residual timing, , of detected photons, along with the angular distribution of hit modules. Equation 4.3 determines the likelihood of observing a charge in a module at a radial distance from the vertex and an opening angle relative to the shower axis , based on an expected charge template (Fig. 7). Finally, the combined likelihood is calculated by simultaneously taking into account the observed photon directions, residual timing, angular distribution of hit modules, and charge in a module (Eq. 4.4). The likelihood product is calculated over per-module detected photons Ns (except Eq. 4.3) and then over hit modules N, and the hypotheses that minimize the negative log-likelihood are identified as the reconstructed parameters. The minimization is achieved with the Nelder-Mead method [Nelder_Mead] with the truth parameter values as the initial seed in order to avoid bias from incomplete minimization. The parameters are illustrated schematically in Fig. 3 where only the scattering effects are shown on the directionality and temporal distributions of observed photons relative to unscattered photons.
@@ 4.3 Photon distributions
The likelihood reconstruction described in Section 4.2 is dependent on the accurate modeling of the underlying PDFs. Due to the grid-like geometry of the detector, the direction, timing, and correlation PDFs, , , , and the expected module charge in a homogeneous ice medium are dependent on the interaction vertex position and direction hypothesis of the shower axis. This is shown in the left panels of Fig. 4 for two combinations of incoming zenith and azimuth of the shower axis for a fixed vertex position and the corresponding angular (top) and temporal (bottom) distributions of detected photons relative to the shower axis. For this study, we shall refer to them as the true integrated PDFs that model the integrated, or cumulative
angular and timing distribution of photons over all hit modules when each event parameter is fixed to the true value.
@@
The features in these PDFs contain information about the incoming directions such as the local geometry around the simulated event in homogeneous ice. Since the scattering process in the ice cannot be solved analytically, these PDFs can only be derived from simulations which means generating new expected distributions for every hypothesis, which soon becomes computationally challenging when reconstructing over a large parameter space. To circumvent the complexity of generating these true PDFs for every hypothesis, we generate representative angular and timing distributions of detected photons using a randomized geometry. The procedure is as follows.
@@
We randomly position 100 simulated strings, each with 50 modules, within a cylindrical volume of base area 40000 m2 and height 200 m. We set thresholds on the minimum and maximum distances between modules to prevent clustering and large gaps. This approach ensures that the instrumentation density of the randomized geometry approximates that of the Upgrade physics region. Events are then simulated with vertices uniformly distributed within a cylindrical region centered at the geometric center of the instrumentation and maintaining a sufficient distance from the edges to prevent photons from escaping. The fitted spatial and temporal distribution of detected photons in all hit modules and averaged over multiple event vertices, referred to as the vertex-averaged integrated PDFs (Fig. 4; middle panel), removes photon sampling bias introduced by the local geometry. They exhibit agreement within the statistical errors (Fig. 4; right panel) for different incoming direction hypotheses due to the homogeneous ice introducing translational and rotational symmetries.
@@
Given that the true integrated PDFs (Fig 4; left panel) differ from averaged integrated PDFs (Fig. 4; middle panel), it is important to assess how this modeling error impacts reconstruction resolutions. We, therefore, evaluate the reconstruction resolutions for two distinct events, with cascade energy of 5 GeV, characterized by different PDFs, one with minimal photon scattering and the other dominated by scattering effects in an Upgrade configuration. The true (averaged) integrated PDFs are shown in blue (green) in Fig. 5 and Fig. 6, with the angular and timing PDFs in the left and middle panels, respectively. We show the 1D projected PDFs without PMT-induced photon direction smearing to highlight the effect of scattering on photon distributions. Note that the subsequent event reconstruction includes all information loss processes, including the photon direction smearing by PMTs. The first case describes a setup where the vertex is 1m away from the nearest module, which is positioned on the Cherenkov ring of the shower. In the second case, the event is directed vertically upwards parallel to a string, with the vertex 15m from the nearest module, resulting in significantly more scattering than in the first scenario.
@@
We define a resolution bias metric as the difference between the zenith-angle resolutions, and , obtained using the averaged and true integrated PDFs, respectively. The results in the right panel of Fig. 5 and Fig. 6 show median differences compatible with zero, indicating that the vertex-averaged PDFs do not yield worse reconstruction on average. The RMS spread, which represents the uncertainty introduced by approximating the PDF, is about 0.06 and 0.09 for the two cases respectively. When compared to the overall resolutions (black line) shown in Fig. 10, with average hit counts of 50 (20) for the less (more) scattering case, the RMS spread is comparable to the median values of 0.1 (0.13). However, the RMS spread is expected to be smaller in less extreme scenarios where the difference in PDFs is less pronounced. It should also be noted here that the true integrated angular and timing distributions sum over photon distributions in all modules and are independent of the source-receiver distance. Due to this, they blur the light pattern information in each module. The improvement in resolutions incorporating the distance-dependent angular and timing distributions is discussed later in Section 5 and presented in Fig. 11.
@@ 4.4 Reconstruction with different observables
Reconstruction is initially performed using the direction (equation 4.1) and residual timing (equation 4.2) information of photon hits to study them individually. The performance metric here is the resolution, quantified as the absolute difference between reconstructed and true cosine zenith. Error bounds are determined as the resolution uncertainty, derived from the 84th and 16th percentiles of the resolution distribution for a given photon count. Since the reconstruction resolution is driven by the number of detected photons, we present the resolutions binned in photon hit counts (instead of cascade energy).
@@
To determine reconstruction resolutions for all cases shown in this study from now on, 30000 purely hadronic events are simulated within a geometry similar to the Upgrade instrumentation, with a uniform module spacing of 20 m horizontally and 3 m vertically. This setup resembles the actual Upgrade geometry and differs from the previously discussed randomized geometry with randomized string and module positions, which is only used to generate the expected PDFs. The events are confined entirely within the detector with shower energies spanning 1-20 GeV. The simulations include all the considered information loss processes, as described in Sections 1.3 and 3. The expected angular and residual timing distribution of hit photons as used in the likelihood PDFs are shown in the middle panel of Fig. 4. The expected charge recorded by a module is modeled by the distribution shown in Fig. 7. Each event is reconstructed using the different likelihood scenarios presented in 4.1, 4.2, 4.3 and then by the combined likelihood 4.4.
@@
The zenith reconstruction performances using the different observables are shown in Fig. 8. Reconstructions that use the direction and timing information of hit photons, shown in dashed green and dotted orange lines, respectively, are comparable suggesting that both are equally significant sources of directional information. The reconstruction based on the charge information from hit modules, represented by the solid magenta line, shows relatively poorer resolution. This outcome is expected since the charge represents the photon summary statistics at each module instead of photon-level information, reducing the available information. The expected distribution p for the combined likelihood is constructed by generating a three-dimensional histogram binned in the angular distribution of hit PMT vectors and hit modules relative to the shower axis and the residual timing for each hit photon. Three-dimensional KDEs are fitted to the underlying distribution and slices showing the marginalized distributions in each dimension are shown in Fig. 12 for proper visualization. We observe that the improvement in median resolutions is minimal compared to those obtained when treating photon direction and timing separately. This suggests that leveraging the correlation between observables does not contribute substantial additional information beyond what is captured by the two-dimensional PDFs discussed in Section 4.3.
@@ 5 Results
We now investigate the processes that limit the physics information in IceCube Upgrade events and degrade reconstruction resolution. Each event is simulated using the same detector setup discussed in Section 3 and with identical MC information before the photon generation and propagation stage. The subsequent simulation incorporates each specific information loss process one at a time, as detailed in Section 3 while keeping all other processes disabled. This simulation step then determines the photon generation, propagation, and detection. Reconstruction is performed using an extended likelihood formulation (Equation 4.4), which offers the best performance as discussed in Section 4.4 and shown in Fig. 8. The expected correlation PDF and module charge, p and p respectively, are generated for each case with the same information loss process included as in the respective simulation.

### v2

@@ 4.1 Reconstruction techniques in IceCube
The longitudinal length of hadronic showers in water and ice increases logarithmically with the shower energy and measures 2-4 meters at energies of O(10 GeV) [Steuer_theses, R_del:2013, shower_KM3NeT]. Since the sparse instrumentation of neutrino telescopes does not allow the resolution of shower details at such low energies, they are modeled as point-like sources with an anisotropic photon emission. A point-like cascade event in IceCube is characterized by 6 parameters [Abbasi:2022]: the incoming zenith angle and azimuth of the shower axis, interaction vertex coordinates = {vx,vy,vz} and the vertex time tv. The average number of emitted photons scales linearly with the summed track length of all charged particles, which increases linearly with the energy of the particle [RADEL2013102, Wiebusch:1995bw]. This results in a sparse pulse distribution, with very few detected photons per event at energies of a few GeV [Abbasi:2022]. Noise cleaning further reduces the physics hits due to misclassification posing challenges to the reconstruction process and necessitating an accurate algorithm.
An IceCube reconstruction algorithm employed in the low-energy regime is the Single-string ANTares-inspired Analysis (SANTA) [Abbasi:2022] which utilizes information from minimally scattered photons. This is done by a rigorous photon hit selection based on the delay timing of photons followed by an event selection based on the surviving photon hit counts. A chi-squared fit on the measured arrival photon hit times for an event hypothesis is finally performed to obtain the optimized parameters. Another standard reconstruction, which additionally exploits the in-ice scattering information from all detected photons, is a likelihood algorithm [Abbasi:2022] known as Reverse Table Reconstruction (RETRO). This method fits a point-like cascade model to the measured light yield in the detector, where the cascade template outlines the expected light distribution in a module for the given hypothesis parameters. A binned Poisson likelihood is then used to determine the hypothesis that best describes the observed light pattern. The likelihood-based approach provides superior resolution in all reconstructed directions but comes at a significantly higher computational cost due to reasons discussed in Section 2.
@@ 4.2 Reconstruction framework for this study
From the discussions in Section 1.2 and 4.1, we model the hadronic cascades as point-like sources in this study. The zenith angle , a key parameter for oscillation analyses, is determined by reconstructing the direction unit vectors of the shower axis. Combined with the parameters listed in the previous section, the study aims to reconstruct 7 parameters - - sufficient to describe the shower topology.
The reconstruction uses an extended unbinned likelihood approach for its statistical robustness, with the assumption that it achieves minimum variance bounds since the assumed PDFs accurately represent the true distributions. Our method uses the direction, residual timing, and charge information of the detected photons, followed by a combined likelihood that exploits the correlation between these observables which is finally used for reconstructions in Section 5. For an anisotropic point source, the angular distribution of the incident photons relative to the direction of the shower axis depends on the shower axis direction because more photons are detected in the modules in front of the shower. For the same reason, the number of photons registered by sensors is direction-dependent because sensors in front of a shower receive more photons than those behind it. The residual timings, defined as the difference between the observed and geometric timestamps, are sensitive to shower direction since sensors behind the shower receive more backscattered and delayed photons.
@@
where is the 7-dimensional parameter set of the event, is the module index, is the photon index for a given module index , and N denotes the total number of modules. The angle enclosed between the shower axis and a detected photon is represented by , and represents the angle between the shower axis and a vector connecting the vertex and a module. The residual timing of a detected photon, defined as the difference between observed and geometric timestamps, is denoted by . The photon direction, timing, and their joint PDFs, , , and are the expected averaged distributions over all hit sensors and are not individually defined for each sensor. The photon count per module, or charge, follows a Poissonian distribution given by
@@
Equation 4.1 calculates the likelihood of observing a given angular distribution of detected photons and hit modules relative to the shower axis (see Fig. 3(a)) for different parameter hypotheses . This formulation exploits the radial symmetry of the photon emission relative to the shower axis to infer its direction. When module resolutions are considered, the photon direction is replaced by the axis direction of the hit PMT. Equation 4.2 evaluates the likelihood based on the observed residual timing of detected photons along with the angular distribution of hit modules (see Fig. 3(b)). Equation 4.3 determines the likelihood of observing a charge in a module at a radial distance from the vertex and an angle relative to the shower axis , based on an expected charge template (Fig. 7). Finally, the combined likelihood is calculated by simultaneously taking into account the observed photon directions, residual timing, angular distribution of hit modules, and charge in a module (Eq. 4.4). The likelihood product is calculated over per-module detected photons Ns (except Eq. 4.3) and then over hit modules N, and the hypotheses that minimize the negative log-likelihood are identified as the reconstructed parameters. The minimization is achieved with the Nelder-Mead method [Nelder_Mead] with the truth parameter values as the initial seed in order to avoid bias from incomplete minimization. The parameters are illustrated schematically in Fig. 3 where only the scattering effects are shown on the directionality and temporal distributions of observed photons relative to unscattered photons.
@@ 4.3 Photon distributions
The likelihood reconstruction described in Section 4.2 is dependent on the accurate modeling of the underlying PDFs. Due to the grid-like geometry of the detector, the direction, timing, and joint PDFs, , , , and the expected module charge in a homogeneous ice medium are dependent on the interaction vertex position and direction hypothesis of the shower axis. This is shown in the left panels of Fig. 4 for two combinations of incoming zenith angle and azimuth of the shower axis for a fixed vertex position and the corresponding angular (top) and temporal (bottom) distributions of detected photons relative to the shower axis. For this study, we shall refer to them as the true integrated PDFs that model the integrated, or cumulative angular and timing distribution of photons over all hit modules when each event parameter is fixed to the true value.
@@
The features in these PDFs contain information about the incoming directions such as the local geometry around the simulated event in homogeneous ice. Since the scattering process in the ice cannot be solved analytically, these PDFs can only be derived from simulations which means generating new expected distributions for every hypothesis, which soon becomes computationally challenging when reconstructing over a large parameter space. To circumvent the complexity of generating these true PDFs for every hypothesis, we generate representative angular and timing distributions of detected photons using a randomized detector geometry. The procedure is as follows.
@@
We randomly position 100 simulated strings, each with 50 modules, within a cylindrical volume of base area 40000 m2 and height 200 m. We set thresholds on the minimum and maximum distances between modules to prevent clustering and large gaps. This approach ensures that the instrumentation density of the randomized geometry approximates that of the Upgrade physics region. The events are then simulated with vertices uniformly distributed within a cylindrical region centered at the geometric center of the instrumentation and maintaining a sufficient distance from the edges to prevent photons from escaping. The fitted spatial and temporal distribution of detected photons across all hit modules and averaged over multiple event vertices are referred to as the vertex-averaged integrated PDFs (Fig. 4; middle panel). The averaging is performed over 10000 events for the 2D histograms (Fig. 4, 5, 6, 7) and over 50000 events for the 3D histograms (Fig. 12) to ensure sufficient MC statistics. The averaging process reduces the photon sampling bias introduced by the local geometry. The integrated PDFs show agreement within the statistical errors (Fig. 4; right panel) for different incoming direction hypotheses due to the homogeneous ice introducing translational and rotational symmetries.
To obtain analytically smooth and continuous functions, variable-bandwidth Kernel Density Estimates (KDE) [VBW_KDE] are fitted to all PDFs in this study. A window moves over all data points and the bandwidth of a Gaussian kernel is calculated depending on the population enclosed within the window. The bandwidth is defined as 1/(population)n to effectively capture the high-statistics regions while smoothing low-statistics areas. Optimization of the bandwidth and window size parameter involves a Kolmogorov-Smirnov goodness-of-fit test [KS_test] which probes parameter values for window size and bandwidth to minimize differences between histogram bin populations and KDE estimates.
@@
Given that the true integrated PDFs (Fig 4; left panel) differ from averaged integrated PDFs (Fig. 4; middle panel), it is important to assess how this modeling error impacts reconstruction resolutions. We therefore evaluate the reconstruction resolutions for two extreme event cases, with cascade energy of 5 GeV, characterized by different PDFs, one with minimal photon scattering and the other dominated by scattering effects in an Upgrade configuration. The true (averaged) integrated PDFs are shown in blue (green) in Fig. 5 and Fig. 6, with the angular and timing PDFs in the left and middle panels, respectively. We show the 1D projected PDFs without PMT-induced photon direction smearing to highlight the effect of scattering on photon distributions. Note that the subsequent event reconstruction includes all information loss processes, including the photon direction smearing by PMTs. The first case describes a setup where the vertex is 1m away from the nearest module, which is positioned on the Cherenkov ring of the shower. In the second case, the event is directed vertically upwards parallel to a string, with its vertex 15m from the nearest module, resulting in significantly more scattering than in the first scenario.
@@
We define a resolution bias metric as the difference between the zenith-angle resolutions, and , obtained using the averaged and true integrated PDFs, respectively. Each of the two extreme event scenarios was simulated 10000 times to construct the PDFs while the reconstruction was performed on 20000 independently simulated events. The results in the right panel of Fig. 5 and Fig. 6 show median differences compatible with zero, suggesting that the vertex-averaged PDFs do not introduce an additional bias on average. The RMS spread, which represents the uncertainty introduced by approximating the PDF, is about 0.06 and 0.09 for the two cases respectively. When compared to the overall resolutions (black line) shown in Fig. 10, with average hit counts of 50 (20) for the less (more) scattering case, the RMS spread is comparable to the median values of 0.1 (0.13). However, the RMS spread is expected to be smaller in less extreme scenarios where the difference in PDFs is less pronounced. The true integrated angular and timing distributions sum over photon distributions in all modules and are independent of the source-receiver distance. Due to this, they blur the light pattern information in each module. The improvement in resolutions incorporating the distance-dependent angular and timing distributions is discussed later in Section 5 and presented in Fig. 11.
@@ 4.4 Reconstruction with different observables
Reconstruction is initially performed using the direction (equation 4.1) and residual timing (equation 4.2) information of photon hits to study them individually. The performance metric here is the resolution, quantified as the absolute difference between reconstructed and true cosine zenith. In NC events, the true direction corresponds to the direction of the hadronic shower axis. In CC and CC events, it lies between the hadronic and electromagnetic shower directions due to momentum transfer between the two shower components [inelasticity]. Error bounds are determined as the resolution uncertainty, derived from the 84th and 16th percentiles of the distribution of residuals for a given photon count. Since the reconstruction resolution is driven by the number of detected photons, we present the resolutions as a function of photon hit counts (instead of cascade energy).
@@
To determine the reconstruction resolutions, 30000 purely hadronic events are simulated for each case within a geometry similar to the Upgrade instrumentation, with a uniform module spacing of 20 m horizontally and 3 m vertically. This setup resembles the Upgrade geometry and differs from the previously discussed randomized geometry with randomized string and module positions, which is only used to generate the expected PDFs. The events are confined entirely within the detector with shower energies uniformly distributed between 1-20 GeV with 1500 events per energy bin. The simulations include all the considered information loss processes, as described in Sections 1.3 and 3. The expected angular and residual timing distribution of hit photons as used in the likelihood PDFs are shown in the middle panel of Fig. 4. The expected charge recorded by a module is modeled by the distribution shown in Fig. 7. Each event is reconstructed using the different likelihood scenarios presented in 4.1, 4.2, 4.3 and then by the combined likelihood 4.4.
@@
The zenith angle reconstruction performances using the different observables are shown in Fig. 8. Reconstructions that use the direction and timing information of hit photons, shown in dashed-dotted green and dashed brown lines, respectively, are comparable suggesting that both are equally significant sources of directional information. The reconstruction based on the charge information from hit modules, represented by the dotted magenta line, shows relatively poorer resolution. This outcome is expected since the charge represents the photon summary statistics at each module instead of photon-level information, reducing the available information. The expected distribution p for the combined likelihood is constructed by generating a three-dimensional histogram as a function of the angular distribution of hit PMT vectors and hit modules relative to the shower axis and the residual timing for each hit photon. Three-dimensional KDEs are fitted to the underlying distribution and slices showing the marginalized distributions in each dimension are shown in Fig. 12 for proper visualization. We observe that the improvement in median resolutions, shown by the solid black line, is minimal compared to those obtained when treating photon direction and timing separately. This suggests that leveraging the correlation between observables does not contribute substantial additional information beyond what is captured by the two-dimensional PDFs discussed in Section 4.3.
@@ 5 Results
We now investigate the processes that limit the physics information in IceCube Upgrade events and degrade reconstruction resolution. Each event is simulated using the same detector setup discussed in Section 3 and with identical MC information before the photon generation and propagation stage. The subsequent simulation incorporates (or ignores) a specific information loss process one at a time, as detailed in Section 3 while keeping all other processes disabled (or enabled). This simulation step then determines the photon generation, propagation, and detection. Reconstruction is performed using an extended likelihood formulation (Equation 4.4), which offers the best performance as discussed in Section 4.4 and shown in Fig. 8. The expected correlation PDF and module charge, p and p respectively, are generated for each case with the same information loss processes included as in the respective simulation.
5.1 Limitations in resolutions from individual information loss processes